Consider the reaction: CO(g)+H2O(g)⇌CO2(g)+H2(g)

K=0.118 at 4000 K

A reaction mixture initially contains a CO partial pressure of 1361 mbar and a H2O partial pressure of 1781 mbar at 4000 K.

Calculate the equilibrium partial pressure of CO

Calculate the equilibrium partial pressure of H2O

Calculate the equilibrium partial pressure of CO2

Calculate the equilibrium partial pressure of H2
Answer #1

Kp is the ratio of partial pressure of products and partial pressure of reactants at equilibrium.

со2 (1) + Ня) са) + нон t=o 38 | һь 1780 мың t-t 136) -Р 1787 -Р кр = •) 8 P.P Ког ». . - P. 362-Р) (1787-Р) Р2. 08 — 242339-P= - 370.756 I J1146556.346 _ ~3700756 I 1070774 1.764 1.764 P=-370.756+ 1070.774 1.764 p:-370.756 – 1070.774 1764 p - P= 36.